# Computation of attractor dimension and maximal sums of Lyapunov exponents using polynomial optimization

### v1

@@ 2.3 Shifted spectrum approach
In the preceding subsection, the possibly unbounded $k$-vectors $y^{k}$ in the tangent space are captured by $z^{k}$, which is bounded because it is the projection of $y^{k}$ onto the unit sphere. In the alternative approach of the present subsection, we capture the dynamics of $y^{k}$ by letting
$w^{k}(t)=w^{k}_{0}e^{-Bt}y^{k}(t),$ (54)
where the initial condition $w^{k}_{0}$ can be any $k$-vector. This $w^{k}(t)$ is bounded forward in time for sufficiently large $B$. The ODE for $w^{k}$ is the same as the ODE 15 for $y^{k}$, except the spectrum of the Jacobian is shifted by $B$. Together with the ODE 6 for $x(t)$ this gives the system
@@
The following lemma asserts that if $B$ is large enough to prevent unbounded growth of $w^{k}$, then $B$ is an upper bound on the LE sum $M_{k}$, and the infimum over such $B$ is equal to $M_{k}$.
@@
Let $\frac{\mathrm{d}}{\mathrm{d}t}x(t)=f(x(t))$ with $f\in C^{1}(\mathcal{B},\mathbb{R}^{n})$ and dynamics forward invariant on $\mathcal{B}\subset\mathbb{R}^{n}$. Let the initial condition $x_{0}$ give a trajectory that is bounded for $t\geq 0$. Let $w^{k}(t)$ satisfy the ODE 55 with initial condition $w_{0}^{k}$. For each $k\in\{1,\ldots,n\}$:
@@
    $M_{k}(x_{0})=\inf_{B\in\mathbb{R}}B\quad\text{s.t.}\quad\text{for each }w^{k}_{0}\in\mathbb{R}^{n_{k}},~w^{k}(t)\text{ is bounded for all }t\geq 0.$ (57)
@@
    For any $B>M_{k}(x_{0})$ there exist $K(x_{0},B)>0$ and $\lambda(x_{0},B)>0$ such that
    $|w^{k}(t)|\leq Ke^{-\lambda t}|w^{k}_{0}|\quad\text{for all $t\geq 0$ and }w^{k}_{0}\in\mathbb{R}^{n_{k}}.$ (58)
  3. 3.
    If $\mathcal{B}$ is compact and $B>\sup_{x_{0}\in\mathcal{B}}M_{k}(x_{0})$, there exist $K(B)$ and $\lambda(B)$ independent of $x_{0}$ such that 58 holds for all $x_{0}\in\mathcal{B}$.
@@
If there exists $y^{k}_{0}$ such that $|y^{k}(t)|\to\infty$ in finite time then, according to 54, no $B$ satisfies the condition in 57. Then both sides of 57 are $\infty$, and the second part holds vacuously. Henceforth we assume that $w(t)$ exists for all $t\geq 0$.
@@
We will first show that if $B$ satisfies the condition of 57 then $M_{k}(x_{0})\leq B$. Substituting $y^{k}=e^{Bt}w^{k}$ into the time average on the right-hand side of 16 gives
$\limsup_{t\to\infty}\frac{1}{t}\log\left|y^{k}(t)\right|$ $=\limsup_{t\to\infty}\frac{1}{t}\log\left(\left|e^{Bt}w^{k}(t)\right|/|w^{k}_{0}|\right)$ (59)
$=B+\limsup_{t\to\infty}\frac{1}{t}\log\left|w^{k}(t)\right|$
$\leq B,$
where the final inequality follows from the boundedness of $w^{k}(t)$. Taking the supremum of the left-hand side over $y^{k}_{0}$ gives $M_{k}(x_{0})\leq B$.
Next we let $B>M_{k}(x_{0})$. Showing that 58 holds will prove the second part of the lemma, and it will complete the proof of the first part by ensuring boundedness of $w^{k}(t)$. Without loss of generality, we assume $|w^{k}_{0}|=1$. Since
$M_{k}(x_{0})$ $\geq\limsup_{t\to\infty}\frac{1}{t}\log{|y^{k}(t)|}$ (60)
$=B+\limsup_{t\to\infty}\frac{1}{t}\log{|w^{k}(t)|},$
we can choose some $\lambda<(B-M_{k}(x_{0}))$ so that $\limsup_{t\to\infty}\frac{1}{t}\log{|w^{k}(t)|}<-\lambda$. Therefore by Grönwall’s inequality there exists a minimal $T(w^{k}_{0})\geq 0$ such that $|w^{k}(t)|\leq e^{-\lambda t}$ for all $t\geq T(w^{k}_{0})$. By compactness, we can let
$T^{*}=\max_{w^{k}_{0}\in\mathbb{S}^{n_{k}-1}}T(w^{k}_{0}),$ (61)
so that $|w^{k}(t)|\leq e^{-\lambda t}$ for all $t\geq T^{*}$ and all $w^{k}_{0}\in\mathbb{S}^{n_{k}-1}$. Finally we let
$K=\max_{\begin{subarray}{c}t\in[0,T^{*}]\\ w^{k}_{0}\in\mathbb{S}^{n_{k}-1}\hskip 1.0pt\end{subarray}}e^{\lambda t}|w^{k}(t)|.$ (62)
which completes the proof of the first two parts of the lemma.
For the third part, assume that $M_{k}(x_{0})$ is bounded above (otherwise the result is vacuously true), so that we can choose $\lambda<B-\sup_{x_{0}\in\mathcal{B}}M_{k}(x_{0})$. Let
$K=\max_{\begin{subarray}{c}x_{0}\in\mathcal{B}\\ t\in[0,T^{\dagger}]\\ w^{k}_{0}\in\mathbb{S}^{n_{k}-1}\hskip 1.0pt\end{subarray}}e^{\lambda t}|w^{k}(t)|,$ (63)
where $T^{\dagger}=\max_{x_{0}\in\mathcal{B}}T^{*}(x_{0})$ for the $T^{*}(x_{0})$ defined in 61. This maximum is valid because $(t,x_{0},w^{k}_{0})\mapsto e^{\lambda t}|w^{k}(t)|$ is continuous on the compact set $0,T^{\dagger}]\times\mathcal{B}\times\mathbb{S}^{n_{k}-1}$. With this value of $\lambda$ and $K$, we must then have 58 for every $x_{0}$, which completes the proof. ∎
@@
Practical upper bounds on the LE sum $M_{k}$ are found by combining 2.10 with the specialised Lyapunov theorem given as B.1 applied to the system 55. This yields the following proposition.
@@
Let $\frac{\mathrm{d}}{\mathrm{d}t}x(t)=f(x(t))$ with $f\in C^{1}(\mathcal{B},\mathbb{R}^{n})$ and dynamics forward invariant on $\mathcal{B}\subset\mathbb{R}^{n}$. Let $m^{B}_{k}(x,w^{k})$ be defined as in 56.
@@
    $\sup_{x_{0}\in\mathcal{B}}M_{k}(x_{0})\leq\inf_{V\in C^{1}(\mathcal{B}\times\mathbb{R}^{n_{k}})}B\quad\text{s.t.}\quad\begin{array}[t]{rl}V\geq&\hskip-7.0pt\left|w^{k}\right|^{2}\\ f\cdot D_{x}V+m^{B}_{k}\cdot D_{w^{k}}V\leq&\hskip-7.0pt0,\end{array}$ (64)
    where the right-hand constraints are imposed for all $(x,w^{k})\in\mathcal{B}\times\mathbb{S}^{n_{k}}$.
@@
    Suppose $\mathcal{B}$ is compact and $f\in C^{2}(\mathcal{B},\mathbb{R}^{n})$. Then 64 holds with equality between the supremum and the infimum, and the Lyapunov functions may be restricted to the form $V(x,w^{k})=(w^{k})^{\mathsf{T}}U(x)w^{k}$ with $U\in C^{1}(\mathcal{B},\mathbb{R}^{n_{k}\times n_{k}})$.
@@
For the first part, assume $B$ and $V$ satisfy the constraints on the right-hand side of 64. Letting $A(x)=Df^{[k]}(x)-BI_{n_{k}}$, the first part of B.1 applies to the system 55 and guarantees boundedness of $w^{k}(t)$ for $t\geq 0$. Boundedness of $w^{k}(t)$ means that $B$ satisfies the constraint on the right-hand side of 2.10, so $M_{k}(x_{0})\leq B$ for all $x_{0}\in\mathcal{B}$. This implies 64, so the first part of the proposition is proven.
@@
For the second part, let $B>\sup_{x_{0}}M_{k}(x_{0})$. By the third part of 2.10, there exist $K>0$ and $\lambda>0$ such that 135 holds for all $x_{0}\in\mathcal{B}$. Observe also that $A\in C^{1}(\mathcal{B},\mathbb{R}^{n_{k}\times n_{k}})$ since $f\in C^{2}$, and that $A(x)$ is bounded uniformly in $x$ since $\mathcal{B}$ is compact. Thus the second part of B.1 applies, so there exists $V(x,w^{k})=(w^{k})^{\mathsf{T}}U(x)w^{k}$ satisfying the constraint in 64. This is true for any $B>\sup_{x_{0}}M_{k}(x_{0})$, so 64 is an equality. ∎
This method gives bounds for the different sums of LEs, which can then be used with proposition 2.7 to bound the Lyapunov dimension. Analogously to proposition 2.8, a direct bound for the dimension is also possible with a modified shifted spectrum approach, using the augmented system
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}x\\ y^{k+1}\\ w\end{bmatrix}=\begin{bmatrix}f(x)\\ Df^{[k+1]}(x)y^{k+1}\\ S_{k}^{B}(x,y^{k+1},w)\end{bmatrix}$, (65)
where we have defined
$S_{k}^{B}(x,y^{k+1},w)=Df^{[k]}(x)w+\frac{B}{1-B}\frac{(y^{k+1})^{\mathsf{T}}Df^{[k+1]}(x)y^{k+1}}{|y^{k+1}|^{2}}w.$ (66)
No shifting is needed in the $y^{k+1}$ dynamics because this quantity decays exponentially on average when $M_{k+1}<0$. The $w$ dynamics 66 shift the $Df^{[k]}$ spectrum not by a negative constant, as in 56, but by a factor proportional to the instantaneous growth rate of $y^{k+1}$, which is negative on average.
@@
Let $\frac{\mathrm{d}}{\mathrm{d}t}x(t)=f(x(t))$ with $f\in C^{1}(\mathcal{B},\mathbb{R}^{n})$ and dynamics forward invariant on $\mathcal{B}\subset\mathbb{R}^{n}$ with each trajectory $x(t)$ bounded for $t\geq 0$. Let the positive integer $k\leq n$ be such that $\sup_{x_{0}\in\mathcal{B}}M_{k+1}(x_{0})<0$. Then the global Lyapunov dimension $d_{L}$ over $\mathcal{B}$ is bounded above by
$d_{L}\leq\inf_{V\in C^{1}}(k+B)\quad\text{s.t.}\quad\begin{array}[t]{rl}V(x,y^{k+1},w)\geq&\hskip-7.0pt\left|w\right|^{2}\\ f\cdot D_{x}V+Df^{[k+1]}y^{k+1}\cdot D_{y^{k}}V+S_{k}^{B}\cdot D_{w}V\leq&\hskip-7.0pt0,\end{array}$ (67)
where the right-hand constraints are imposed for all $(x,y^{k+1},w)\in\mathcal{B}\times\mathbb{R}^{n_{k+1}}\times\mathbb{R}^{n_{k}}$, and $D_{x}V$, $D_{z^{k}}V$ and $D_{z^{k+1}}V$ denote the gradients of $V(x,y^{k+1},w)$ with respect to each argument.
@@
We will prove that, for any $B$ and $V$ satisfying the constraints in 67, $d_{L}\leq k+B$. We will use 2.5 to do this. First note that $\sup_{x_{0}\in\mathcal{B}}M_{k+1}(x_{0})<0$ implies that $k\geq j$ where $j$ is the minimal integer of 2.3. Observe that the dynamics of $w$ in the system 65 are of the form $\frac{\mathrm{d}w}{\mathrm{d}t}=A(x,y^{k+1})w$ for some matrix $A$, so we may apply B.1 to this system. This tells us that given any $B$ and $V$ satisfying the constraints, $w(t)$ is bounded for any initial conditions. We now suppose that we have such a $V$ and $B$ and show that 23 holds for all $x_{0}$.
Consider any initial conditions for 65. Let $y^{k}=|y^{k+1}|^{\frac{-B}{1-B}}w$, which combined with $6$ implies that $y^{k}$ satisfies the usual tangent dynamics 15. We then have
$\frac{w^{\mathsf{T}}Df^{[k]}w}{|w|^{2}}=\frac{(y^{k})^{\mathsf{T}}Df^{[k]}y^{k}}{|y^{k}|^{2}}.$ (68)
Since $|w(t)|$ is bounded,
$0$ $\geq\limsup_{T\to\infty}\frac{1}{T}\log\frac{|w(T)|}{|w(0)|}$ (69)
$=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}\frac{\mathrm{d}}{\mathrm{d}t}\log\frac{|w(t)|}{|w(0)|}\mathrm{d}t$ (70)
$=\limsup_{T\to\infty}\frac{1}{T}\int_{0}^{T}\frac{1}{2}\left(\frac{\frac{\mathrm{d}|w|^{2}}{\mathrm{d}t}}{|w|^{2}}\right)\mathrm{d}t$ (71)
$=\limsup_{T\to\infty}\frac{1}{T}\int_{0}^{T}\left(\frac{w^{\mathsf{T}}(Df)^{[k]}w}{|w|^{2}}+\frac{B}{1-B}\frac{(y^{k+1})^{\mathsf{T}}(Df)^{[k+1]}y^{k+1}}{|y^{k+1}|^{2}}\right)\mathrm{d}t.$ (72)
Using 68 and 26, this gives
$0\leq\limsup_{t\to\infty}\frac{1}{t}\log{|y^{k}(t)|}+\frac{B}{1-B}\limsup_{t\to\infty}\frac{1}{t}\log{|y^{k+1}(t)|}.$ (73)
Since this holds for all $y^{k+1}$ and $y^{k}$, by 27 we must have
$0\leq M_{k}(x_{0})+\frac{B}{1-B}M_{k+1}(x_{0}).$ (74)
which rearranges to
$B(M_{k}(x_{0})-M_{k+1}(x_{0}))-M_{k}(x_{0})\geq 0.$ (75)
2.5 finishes the proof. ∎

### v2

@@ 2.3 Shifted spectrum approach
In the preceding subsection, the possibly unbounded $k$-vectors $y^{k}$ are replaced by their projections $z^{k}$ onto the unit sphere, which are necessarily bounded. In the alternative approach of the present subsection, we replace the dynamics of $y^{k}$ by
$w^{k}(t)=c\hskip 1.0pte^{-Bt}y^{k}(t)$ (54)
for any nonzero constant $c$, so the initial condition $w_{0}^{k}=c\hskip 1.0pty_{0}^{k}$ can be any nonzero point in $\mathbb{R}^{C_{n}^{k}}$. This $w^{k}(t)$ is bounded forward in time for sufficiently large $B$. The ODE for $w^{k}$ is the same as the ODE 15 for $y^{k}$, except the spectrum of the Jacobian is shifted by B-B. Together with the ODE 6 for $x$ this gives the augmented system
@@
Note that this ODE is linear in $w^{k}$, whereas the ODE 29 in the sphere projection approach is nonlinear in $z^{k}$. The following lemma asserts that if $B$ is large enough to prevent unbounded growth of $w^{k}$, then $B$ is an upper bound on the LE sum $M_{k}$, and the infimum over such $B$ is equal to $M_{k}$. The second part gives stronger conditions for this to be true in a uniform sense for all initial conditions.
@@
Let $x(t)$ evolve by the ODE 6, where $f\in C^{1}(\mathcal{B},\mathbb{R}^{n})$ and $\mathcal{B}$ is forward invariant, and let $w^{k}(t)$ evolve by the ODE 55 with $B$ fixed. Assume each trajectory $x(t)$ in $\mathcal{B}$ is bounded for $t\geq 0$. For each $k\in\{1,\ldots,n\}$:
@@
    $M_{k}(x_{0})=\inf_{B\in\mathbb{R}}B\quad\text{s.t.}\quad\text{for each }w^{k}_{0}\in\mathbb{R}^{C_{n}^{k}},~w^{k}(t)\text{ is bounded for }t\geq 0.$ (57)
@@
    Suppose $f\in C^{2}(\mathcal{B},\mathbb{R}^{n})$. For each compact set $\mathcal{A}\subset\mathcal{B}$ and each $B>\sup_{x_{0}\in\mathcal{A}}M_{k}(x_{0})$, there exist constants $K>0$ and $\lambda>0$ such that
    $|w^{k}(t)|\leq Ke^{-\lambda t}|w^{k}_{0}|\quad\text{for all $t\geq 0$ and }(x_{0},w^{k}_{0})\in\mathcal{A}\times\mathbb{R}^{C_{n}^{k}}.$ (58)
@@
We first show that if $B$ satisfies the constraint in 57, then $M_{k}(x_{0})\leq B$. Substituting $y^{k}=e^{Bt}w^{k}/c$ into the time average on the right-hand side of 16 gives
$\limsup_{t\to\infty}\frac{1}{t}\log\left|y^{k}(t)\right|=\limsup_{t\to\infty}\frac{1}{t}\log\left(\left|e^{Bt}w^{k}(t)\right|/c\right)=B+\limsup_{t\to\infty}\frac{1}{t}\log\left|w^{k}(t)\right|\leq B,$ (59)
where the final inequality follows from boundedness of $w^{k}(t)$. Taking the supremum of the left-hand side over $y^{k}_{0}$ gives $M_{k}(x_{0})\leq B$. Next, we assume that $B>M_{k}(x_{0})$ and show that the constraint in 57 must hold. With such $B$, and with $\mathcal{A}=\{x_{0}\}$, the second part of the proposition (proved below) applies because $B>M_{k}(x_{0})=\sup_{x_{0}^{\prime}\in\mathcal{A}}M_{k}(x_{0}^{\prime})$. Thus 58 holds and implies that $w^{k}(t)$ is bounded for all $t\geq 0$, and the first part is proved.
To prove the second part of the proposition, let $B>\sup_{x_{0}\in\mathcal{A}}M_{k}(x_{0})$. It suffices to show that $K$ and $\lambda$ exists so that 58 holds for $(x_{0},w_{0}^{k})\in\mathcal{A}\times\mathbb{S}^{C_{n}^{k}-1}$, rather than all $w^{k}_{0}\in\mathbb{R}^{C_{n}^{k}-1}$. These same $K$ and $\lambda$ remain valid for all $w^{k}_{0}\in\mathbb{R}^{C_{n}^{k}}$ since $w^{k}(t)$ scales proportionally to $w^{k}_{0}$, due to the linearity of the $w^{k}$ dynamics in 55.
To choose $\lambda$, note that
$\sup_{x_{0}\in\mathcal{A}}M_{k}(x_{0})$ $=\sup_{(x_{0},y^{k}_{0})\in\mathcal{A}\times\mathbb{S}^{C_{n}^{k}-1}}\limsup_{t\to\infty}\frac{1}{t}\log{|y^{k}(t)|}$ (60)
$=B+\sup_{(x_{0},w^{k}_{0})\in\mathcal{A}\times\mathbb{R}^{C_{n}^{k}}}\limsup_{t\to\infty}\frac{1}{t}\log{|w^{k}(t)|},$
and so
$\sup_{(x_{0},w^{k}_{0})\in\mathcal{A}\times\mathbb{R}^{C_{n}^{k}}}\limsup_{t\to\infty}\frac{1}{t}\log{|w^{k}(t)|}<0.$ (61)
We therefore can choose a $\lambda>0$ such that
$\limsup_{t\to\infty}\frac{1}{t}\log{|w^{k}(t)|}<-\lambda\quad\text{ for all }\quad(x_{0},w^{k}_{0})\in\mathcal{A}\times\mathbb{R}^{C_{n}^{k}}.$ (62)
To choose $K$, note that 62 implies $|w^{k}(t)|\leq e^{-\lambda t}|w^{k}_{0}|$ after some time $T(x_{0},w^{k}_{0})$. We choose $T$ to depend continuously on the initial conditions $(x_{0},w^{k}_{0})$, which is possible since continuity of $w^{k}(t)$ in both $t$ and $(x_{0},w^{k}_{0})$ follows from the fact that the right-hand side of the ODE 55 is locally Lipschitz when $f\in C^{2}$. Having chosen $T(x_{0},w^{k}_{0})$ to be continuous, the maximum
$T^{*}=\max_{(x_{0},w^{k}_{0})\in\mathcal{A}\times\mathbb{S}^{C_{n}^{k}-1}}T(x_{0},w^{k}_{0})$ (63)
is finite since it is over a compact set. Therefore $|w^{k}(t)|\leq e^{-\lambda t}|w^{k}_{0}|$ for all $t>T^{*}$ and all $(x_{0},w_{0})\in\mathcal{A}\times\mathbb{S}^{C_{n}^{k}-1}$. Finally, we choose
$K=\max_{\begin{subarray}{c}(x_{0},w^{k}_{0})\in\mathcal{A}\times\mathbb{S}^{C_{n}^{k}-1}\\ t\in[0,T^{*}]\end{subarray}}e^{\lambda t}|w^{k}(t)|,$ (64)
which again is finite since it is the maximum of a continuous function over a compact set. For the chosen $K$ and $\lambda$, the inequality in 58 indeed holds over $\mathcal{A}\times\mathbb{S}^{C_{n}^{k}-1}$ and thus also over $\mathcal{A}\times\mathbb{R}^{C_{n}^{k}}$. ∎
@@
Practical upper bounds on the LE sum $M_{k}$ are found by applying a specialized function theorem (B.1) to the ODE system 55 and combining the result with 2.10. This yields the following proposition.
@@
Let $x(t)$ evolve by the ODE 6, where $f\in C^{1}(\mathcal{B},\mathbb{R}^{n})$ and $\mathcal{B}$ is forward invariant. Let $m^{B}_{k}(x,w^{k})$ be as defined in 56.
@@
    Suppose each trajectory $x(t)$ in $\mathcal{B}$ is bounded for $t\geq 0$. For each $k\in\{1,\ldots,n\}$, the maximum leading LE sum $M_{k}$ is bounded above by
    $\sup_{x_{0}\in\mathcal{B}}M_{k}(x_{0})\leq\inf_{V\in C^{1}(\mathcal{B}\times\mathbb{R}^{C_{n}^{k}})}B\quad\text{s.t.}\quad\begin{array}[t]{rl}V\geq&\hskip-7.0pt\left|w^{k}\right|^{2}\\ f\cdot D_{x}V+m^{B}_{k}\cdot D_{w^{k}}V\leq&\hskip-7.0pt0,\end{array}$ (65)
    where the right-hand constraints are imposed for all $(x,w^{k})\in\mathcal{B}\times\mathbb{R}^{C_{n}^{k}}$.
@@
    Suppose $\mathcal{B}$ is compact and $f\in C^{2}(\mathcal{B},\mathbb{R}^{n})$. Then 65 holds with equality between the supremum and the infimum, including when Lyapunov functions are restricted to the form $V(x,w^{k})=(w^{k})^{\mathsf{T}}U(x)w^{k}$ with $U\in C^{1}(\mathcal{B},\mathbb{R}^{C_{n}^{k}\times C_{n}^{k}})$.
@@
For the first part, assume $B$ and $V$ satisfy the constraints on the right-hand side of 65. Letting $A(x)=Df^{[k]}(x)-BI_{C_{n}^{k}}$, the first part of B.1 applies to the system 55 and guarantees the boundedness of $w^{k}(t)$ for $t\geq 0$. Boundedness of $w^{k}(t)$ means that $B$ satisfies the constraint on the right-hand side of 57, so $M_{k}(x_{0})\leq B$ for all $x_{0}\in\mathcal{B}$. This implies 65, so the first part of the proposition is proved.
@@
For the second part, let $B>\sup_{x_{0}\in\mathcal{B}}M_{k}(x_{0})$. By the second part of 2.10, there exist $K,\lambda>0$ such that 126 holds. Since $A\in C^{1}(\mathcal{B},\mathbb{R}^{C_{n}^{k}\times C_{n}^{k}})$ when $f\in C^{2}$, and since $\mathcal{B}$ is compact, the second part of B.1 guarantees the existence of $V(x,w^{k})=(w^{k})^{\mathsf{T}}U(x)w^{k}$ satisfying the constraint in 65. This is true for any $B>\sup_{x_{0}\in\mathcal{B}}M_{k}(x_{0})$, so 65 is an equality. ∎
To bound the Lyapunov dimension $d_{L}$ using the shifted spectrum approach, one option is to find a negative upper bound on $M_{k+1}$ and a nonnegative upper bound on $M_{k}$, and then use expression 33 of proposition 2.7. Another option, which is potentially sharper, is to bound $d_{L}$ directly, similarly to proposition 2.8 in the sphere projection approach. However, whereas proposition 2.8 avoids $y^{k}$ and $y^{k+1}$ by using their projections onto the unit sphere, here we retain the $y^{k+1}$ variable and shift the spectrum of $y^{k}$ by a factor that depends on $y^{k+1}$. We assume that $k$ is chosen so that $d_{L}<k+1$, as can be confirmed by a negative upper bound on $M_{k+1}$. Then $y^{k+1}(t)$ is bounded and decays exponentially on average. The variable $y^{k}(t)$ might be unbounded, so we instead use
$w(t)=\big|y^{k+1}(t)\big|^{\tfrac{B}{1-B}}y^{k}(t),$ (66)
which will be bounded for large enough $B\in[0,1)$ since $y^{k+1}(t)$ decays. The particular expression 66 for $w$ is chosen so that boundedness of $w$ implies $d_{L}\leq k+B$, as asserted by 2.12 below. (We do not pursue conditions under which the the bound on $d_{L}$ given by this lemma is an equality.)
Lemma 2.12.
Let $x(t)$ evolve by the ODE 6, where $f\in C^{1}(\mathcal{B},\mathbb{R}^{n})$ and $\mathcal{B}$ is forward invariant. Assume each trajectory $x(t)$ in $\mathcal{B}$ is bounded for $t\geq 0$. For any positive integer $k<n$ such that $\sup_{x_{0}\in\mathcal{B}}M_{k+1}(x_{0})<0$, let $w(t)$ be as defined by $6$, where $y^{k}(t)$ and $y^{k+1}(t)$ evolve by 15. Then the global Lyapunov dimension $d_{L}$ over $\mathcal{B}$ is bounded above by
$d_{L}\leq\inf_{B\in[0,1)}(k+B)\quad\text{s.t.}\quad\begin{array}[t]{l}\text{for each }(x_{0},y_{0}^{k},y_{0}^{k+1})\in\mathcal{B}\times\mathbb{R}^{C_{n}^{k}}\times\mathbb{R}^{C_{n}^{k+1}},\\ ~w(t)\text{ is bounded for }t\geq 0.\end{array}$ (67)
Proof.
Let $B\in[0,1)$ be such that $w(t)$ defined by 66 is bounded for $t\geq 0$ with each initial condition. It suffices to show that $d_{L}\leq k+B$. Taking the logarithm of $|w(t)|$ and dividing by $t>0$ gives
$\frac{1}{t}\log|w(t)|=\frac{B}{1-B}\frac{1}{t}\log\big|y^{k+1}(t)\big|+\frac{1}{t}\log\big|y^{k}(t)\big|.$ (68)
The left-hand side is nonpositive as $t\to\infty$ since $w(t)$ is bounded, so taking this limit and then maximizing over initial conditions gives
$0\geq\left(\frac{B}{1-B}\right)\sup_{y^{k+1}_{0}\in\mathbb{S}^{{C_{n}^{k+1}}-1}}\limsup_{t\to\infty}\frac{1}{t}\log\big|y^{k+1}(t)\big|+\sup_{y^{k}_{0}\in\mathbb{S}^{{C_{n}^{k}}-1}}\limsup_{t\to\infty}\frac{1}{t}\log{|y^{k}(t)|}.$ (69)
Recalling expression 16 for $M_{k}$, we see that 69 is equivalent to $B\left[M_{k}(x_{0})-M_{k+1}(x_{0})\right]-M_{k}(x_{0})\geq 0$, and so 2.5 gives the desired conclusion that $d_{L}\leq k+B$ for all $x_{0}$. ∎
In order to apply 2.12 to a particular ODE, one must show that $w(t)$ indeed remains bounded for all initial conditions. Such boundedness can be shown using the specialized Lyapunov function result of B.1, for which we need the dynamics of $w(t)$. An ODE for $w$ is found by taking the time derivative of 66, using the ODE 7 for $y^{k}$ and $y^{k+1}$, and eliminating $y^{k}$ in favor of $w$. The resulting ODE, together with the ODEs for $x$ and $y^{k+1}$, form the augmented system
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}x\\ w\\ y^{k+1}\end{bmatrix}=\begin{bmatrix}f(x)\\ p_{k}^{B}(x,w,y^{k+1})\\ Df^{[k+1]}(x)y^{k+1}\end{bmatrix}$, (70)
where
$p_{k}^{B}(x,w,y^{k+1})=Df^{[k]}(x)w+\frac{B}{1-B}\frac{(y^{k+1})^{\mathsf{T}}Df^{[k+1]}(x)y^{k+1}}{|y^{k+1}|^{2}}w.$ (71)
B.1 can be used to show boundedness of $w(t)$ that satisfy 70 for chosen $k$ and $B$, in which case 2.12 will imply $d_{L}\leq k+B$. The following proposition states this main result for bounding Lyapunov dimension using the shifted spectrum approach.
@@
Let $x(t)$ evolve by the ODE 6, where $f\in C^{1}(\mathcal{B},\mathbb{R}^{n})$ and $\mathcal{B}$ is forward invariant. Assume each trajectory $x(t)$ in $\mathcal{B}$ is bounded for $t\geq 0$. For any positive integer $k<n$ such that $\sup_{x_{0}\in\mathcal{B}}M_{k+1}(x_{0})<0$, let $p_{k}^{B}(x,w,y^{k+1})$ be as defined by 71 . Then the global Lyapunov dimension $d_{L}$ over $\mathcal{B}$ is bounded above by
$d_{L}\leq\inf_{\begin{subarray}{c}B\in[0,1)\\ V\in C^{1}\hskip 6.0pt\end{subarray}}(k+B)\quad\text{s.t.}\quad\begin{array}[t]{rl}V(x,w,y^{k+1})\geq&\hskip-7.0pt\left|w\right|^{2}\\ f\cdot D_{x}V+p_{k}^{B}\cdot D_{w}V+\big(Df^{[k+1]}y^{k+1}\big)\cdot D_{y^{k+1}}V\leq&\hskip-7.0pt0,\end{array}$ (72)
where the right-hand constraints are imposed for all $(x,w,y^{k+1})\in\mathcal{B}\times\mathbb{R}^{C_{n}^{k}}\times\mathbb{R}^{C_{n}^{k+1}}$, and $D_{x}V$, $D_{w}V$ and $D_{y^{k+1}}V$ denote the gradients of $V(x,w,y^{k+1})$ with respect to each argument.
@@
Let $B$ and $V(x,w,y^{k+1})$ satisfy the constraints in 72. It suffices to show that $d_{L}\leq k+B$. The inequality constraints are exactly the Lyapunov function conditions 125 in B.1, written for the case where the nonlinear variables are $(x,y^{k+1})$ and evolve as in 70, and the linear variable $w$ evolves as $\frac{\mathrm{d}}{\mathrm{d}t}w=A(x,y^{k+1})w$ with the form of $A$ implied by 71. Thus, by the first part of B.1, the constraints in 72 imply that $w(t)$ is bounded for all initial conditions $(x_{0},w_{0},y_{0}^{k+1})$. 2.12 then implies that $d_{L}\leq k+B$, which completes the proof. ∎